OpenStudy (anonymous):

Solve, using the substitution method. 4x + 2y = 11 x – 2 = – 2y

OpenStudy (anonymous):

Is \[2y=11-4x\] Right?

OpenStudy (potatoes.ramu):

Multiply the second equation by 2

OpenStudy (potatoes.ramu):

and then subtract it from the first

OpenStudy (anonymous):

Ok, so \[2x-4=-4y?\]

OpenStudy (potatoes.ramu):

Wait.. actually multiply the second equation by 4. That way you eliminate the x term

OpenStudy (anonymous):

\[4x-8=-8y\]

OpenStudy (anonymous):

How would I subtract it from the first equation? X becomes eliminated than what?

OpenStudy (anonymous):

@hartnn

OpenStudy (potatoes.ramu):

When x is eliminated, you can find the y term

hartnn (hartnn):

substitution method. (not elimination method) 4x + 2y = 11 x – 2 = – 2y (find 'x' from here, by adding 2 on both sides)

OpenStudy (potatoes.ramu):

Once you find the y term, you can substitute it in any of the above equations, and then you get x.

OpenStudy (anonymous):

x=y?

OpenStudy (anonymous):

So 6y=11?

OpenStudy (potatoes.ramu):

@hartnn you'll still have the y term

hartnn (hartnn):

x – 2 = – 2y x = 2-2y plug this in 1st equation 4(2-2y) +2y =11 simplify and get y

hartnn (hartnn):

ask if you don't get that ^

OpenStudy (potatoes.ramu):

@hartnn , but can't you solve it using elimination as well?

OpenStudy (anonymous):

8-8y+2y=11

OpenStudy (anonymous):

-10y=3?

hartnn (hartnn):

the system can be solved by elimination method, but since the question specifically asks for substitution, we need to use substitution

hartnn (hartnn):

-8y +2y = ... ?

OpenStudy (anonymous):

6y

OpenStudy (anonymous):

8-6y=11 -6y=3 y=-1/2

hartnn (hartnn):

YES :)

hartnn (hartnn):

now find x

OpenStudy (anonymous):

Okkkkk :)

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

So its this bit.ly/1ewUix3

hartnn (hartnn):

\(\huge \checkmark \)

OpenStudy (anonymous):

(3,-1/2) Right?

hartnn (hartnn):

yes, it is correct.

OpenStudy (anonymous):

Thanks! Should I give the medal to potatoes or you? He tried... but you helped

hartnn (hartnn):

i'm not here for medals. you understand thats my medal :)

OpenStudy (anonymous):

:D

OpenStudy (potatoes.ramu):

Also.. I'm a she :) And my method was wrong for this problem! Sorry!